Aggregator

COMMENT ON MST magnet installation and related magnet work at 1L09,2L03, 2L25, 1L26

6 days 22 hours ago
Log Number: 4467938Logbook(s): Body: 

Mag Shop:

A brief note for today's work. We started working @2L03 and quickly found a major error. When the VAC group finished their work, they had tried to attach LCW unions at magnet mid plane on up and dn beam end of QB quad. When LCW was cut on LCW was flying everywhere. This cost us an entire day to recover. When you break a LCW cooling circuit inside the quad a portion of the magnet has to be removed as the damage is extensive. We located a QB quad at shop and completed QA to ensure it was a good functional quad.

jbeck

Autolog save

6 days 23 hours ago
Log Number: 4467931Logbook(s): Tags: Body: saver ran from opsl632

Saved: samari 01-30-26 16:06:32 {A:5 B:3 C:1 E:340 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in. dump(s) in. Autosave prior to restore.
samari

Vacuum Group Summary 2026-01-30

6 days 23 hours ago
Log Number: 4467929Logbook(s): Body: 

2L25
- VBV2L25B and VBV2L26A replaced
- Restore 2L24 ins. vac. next
- VIP2L27A pumping on module over weekend
- Prep work for C100 module swap continues
- C100 girder assembly in progress in clean room

duzik

Operator SIGN IN

6 days 23 hours ago
Log Number: 4467928Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Anthony
DATE:             01/30/2026
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log
anthony

COMMENT ON RF work for 2026 HCO

6 days 23 hours ago
Log Number: 4467927Logbook(s): Body: 

Several zones in the North and South linacs have Sorenson power supplies that provide current to the C100 solenoids. These power supplies need maintenance every year due to getting hot and shutting down. We currently have six and possibly 7 in the North Linac zones 1L24 and 1L25 that need pulled and cooling fans checked/replaced. These two zones wouldn't go into HV for PSS certification this week and will not be ready until the supplies are repaired and reinstalled so that we can go into HV in the zone. Larry Farrish has 6 in for repair that hopefully will work early next week. Kerns 7283.

kerns

COMMENT ON Investigate Trim Rack Problems Found doing HCO

6 days 23 hours ago
Log Number: 4467925Logbook(s): Body: 

We had to re-connect a comm cable inside the Utility Chassis for E301B11 Trim Rack, all magnets where set to +/- low current for 30 min and then cycled. All magnets and Trim Rack are ready for operation.

We had to replace a neg. bootstrap fuse on the back of the Utility Chassis in W301B13 Trim Rack. all magnets where set to +/- low current for 30 min and then cycled. All magnets and Trim Rack are ready for operation.

We had to replace a front door switch in E501B11 and a back door switch in E501B12 Trim Rack. all magnets where set to +/- low current for 30 min and then cycled.

diggs

COMMENT ON Wien filter chiller relocation and H-Wien cooling plates installation

6 days 23 hours ago
Log Number: 4467924Logbook(s): Body: 

Thermocouple wire terminations have been completed today by Kevin with commercial ends wrapped in Kapton. Kevin installed a connector block where the TC wires coming from the Labjack T7 transition to the TC terminations. The V-Wien TCs are taped on top (1I04 Ch. 1) and bottom (1I04 Ch. 2) of the magnet yoke. The H-Wien TCs are taped on the magnet yoke beam right (1I06 Ch. 1) and beam left (1I06 ch. 2).

PVs for these channels are now in EPICS reporting in Celsius, and are being archived: https://logbooks.jlab.org/entry/4467142

chgarcia

HYST AREA 7-0 Cycles Magnets

1 week ago
Log Number: 4467923Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MQA7S04
MBC9S02H
MBC9S02V
MQA3S04
MBDBS01H
MBDBS01V
MBT1S05H
MBT1S05V
MQB1S06
MQD9S02
MQA9S02
MQC9S00
MBC7S03V
MQA9S01
MBC9S01H
MBC9S01V

Total Time: 1.933 mins

diggs

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Blackshaw
MCC Operators Janak
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph